# A Refreshed Similarity-based Upsampler for Direct High-Ratio Feature Upsampling

### v1

@@ Abstract
Feature upsampling is a fundamental and indispensable ingredient of almost all current network structures for image segmentation tasks. Very recently, a popular similarity-based feature upsampling pipeline has been proposed, which utilizes a high-resolution (HR) feature as guidance to help upsample the low-resolution (LR) deep feature based on their local similarity. Albeit achieving promising performance, this pipeline has specific limitations in methodological designs: 1) HR query and LR key features are not well aligned in a controllable manner; 2) the similarity between query-key features is computed based on the fixed inner product form, lacking flexibility; 3) neighbor selection is coarsely operated on LR features, resulting in mosaic artifacts. These shortcomings make the existing methods along this pipeline primarily applicable to hierarchical network architectures with iterative features as guidance and they are not readily extended to a broader range of structures, especially for a direct high-ratio upsampling. Against these issues, we thoroughly refresh this pipeline and meticulously optimize every methodological design. Specifically, we firstly propose an explicitly controllable query-key feature alignment from both semantic-aware and detail-aware perspectives, and then construct a parameterized paired central difference convolution block for flexibly calculating the similarity between the well-aligned query-key features. Besides, we develop a fine-grained neighbor selection strategy on HR features, which is simple yet effective for alleviating mosaic artifacts. Based on these careful designs, we systematically construct a refreshed similarity-based feature upsampling framework named ReSFU. Comprehensive experiments substantiate that our proposed ReSFU is finely applicable to various types of architectures in a direct high-ratio upsampling manner, and consistently achieves satisfactory performance on different applications, including semantic segmentation, medical image segmentation, instance segmentation, and panoptic segmentation, showing superior generality and ease of deployment beyond the existing upsamplers. Codes are available at https://github.com/zmhhmz/ReSFU.
@@ 1 Introduction
Figure 1: Upper: For the ×$\times 8$ upsampling from $C4$ to $\widetilde{C4}$ in the hierarchical SegFormer [1], (a) SAPA [2] iteratively performs three ×$\times 2$ upsampling processes with intermediate features as guidance; (b) our ReSFU obtains clearer object boundaries in a direct ×$\times 8$ upsampling only with the guidance $C1$. Lower: For the ×$\times 4$ upsampling from $\bm{x}$ to $\tilde{\bm{x}}$ with the direct guidance of the shallow HR feature $\bm{y}$ in the non-hierarchical Segmenter [3], (c) SAPA suffers from mosaic artifacts, while (d) our ReSFU achieves better structural preservation. All the images are best viewed by zooming in on screen, especially to observe mosaic artifacts.
As a fundamental ingredient in deep network architectures, feature upsampling aims to restore the spatial resolution of low-resolution (LR) features, which is widely used in image segmentation tasks, such as semantic segmentation [4], instance segmentation [5], and panoptic segmentation [6].
@@
Against this issue, a novel similarity-based feature upsampling pipeline SAPA [2] has been proposed, which utilizes an HR feature $\bm{y}$ as the guidance to help accomplish the upsampling process from the LR feature $\bm{x}$ to the HR one $\tilde{\bm{x}}$. Taking the representative base version of SAPA as an example, the HR feature element $\tilde{\bm{x}}_{i}$ at pixel $i$ is estimated as (see Sec. 3 for more details):
@@
$\tilde{\bm{x}}_{i}=\underbracket{\text{Softmax}(\bm{q}_{i}\bm{k}^{T}_{\mathcal{N}(i)})}_{\text{upsampling kernel}}\bm{x}_{\mathcal{N}(i)},$ (1)
where the HR query $\bm{q}$ and the LR key $\bm{k}$ are linear projections of $\bm{y}$ and $\bm{x}$, respectively; $\mathcal{N}(i)$ is the neighborhood of $i$; the upsampling kernel is computed on the inner produced-based similarity between the query-key pair. We can easily see that the entire upsampling framework along the similarity-based pipeline is closely related to the design of three main components: ❶ the acquisition of query-key features; ❷ the similarity calculation manner between the query-key pair; ❸ the neighbor selection $\mathcal{N}(i)$.
@@
Albeit obtaining impressive success, we carefully delve into this pipeline and find that each of the aforementioned components in SAPA still suffers from unresolved issues. Specifically, ❶ considering that the HR guidance feature $\bm{y}$ is usually shallower than the LR deep feature $\bm{x}$, only with the content-agnostic linear projection operation, the generated $\bm{q}$ and $\bm{k}$ are generally not well semantically aligned in a controllable manner. This would adversely interfere with the subsequent similarity calculation; ❷ The inner product-based similarity calculation manner in Eq. (1) lacks enough flexibility and often results in blurry effects in the upsampled features (see $\widetilde{C4}$ in Fig. 1 (a)); ❸ From Eq. (1), the neighbor selection $\mathcal{N}(i)$ is coarsely executed on LR features $\bm{k}$ and $\bm{x}$, which would incur mosaic artifacts, especially in a direct high-ratio upsampling (see $\tilde{\bm{x}}$ in Fig. 1 (c)).
@@
These three methodological design limitations weaken the potential application of SAPA, making it usually only applicable to iterative ×$\times 2$ upsampling structures with hierarchical guidances but less suitable for direct high-ratio upsampling without iterative features as guidance. For instance, for the hierarchical architecture SegFormer [1] with four levels of features as shown in Fig. 1 (a), to accomplish the ×$\times 8$ upsampling from $C4$ to $\widetilde{C4}$, SAPA, as well as other existing feature upsampling methods [7, 12, 13], performs three consecutive ×$\times 2$ upsampling processes with $C3,C2$, and $C1$ as guidance features, respectively. Such an iterative pattern definitely leads to the laborious deployment of more upsampling modules. Besides, for the non-hierarchical structures, such as Segmenter [3] in Fig. 1 (c), the deep LR feature $\bm{x}$ requires a direct ×$\times 4$ upsampling under the guidance of a shallow HR feature $\bm{y}$. Without iterative guidance features, the upsampled feature $\tilde{\bm{x}}$ exhibits adverse visual artifacts. These issues generally exist in most of the current feature upsampling methods (see Sec. 5.2). Faced with this tricky scenario, is it possible to build an architecture-agnostic feature upsampling framework that is applicable to different types of network architectures and can always achieve satisfactory effects in direct high-ratio upsampling without complicated iterative processes?
To answer this question, in this paper, we meticulously refresh the pipeline in Eq. (1) and carefully optimize the involved three main components one by one. Then we propose a Refreshed Similarity-based Feature Upsampling framework, called ReSFU, with higher flexibility and better universality, which can be easily embedded in a direct high-ratio upsampling manner for different architectures as shown in Figs. 1 (b) and (d). Specifically, our contributions are mainly four-fold:
1) Firstly, we propose an explicitly controllable algorithm to finely align query-key features from both semantic-aware and detail-aware perspectives. Concretely, we utilize guided filter (GF) [14] to explicitly optimize the original linearly-projected HR query $\bm{q}$ and execute a controllable transformation on it to promote semantic alignment with the key. Meanwhile, we further investigate the HR query $\bm{q}$ itself and adopt an explicit Gaussian smoothing to help accomplish the HR feature self-alignment for better detail preservation. Such careful query-key alignment designs facilitate more accurate similarity computation and make it possible to flexibly select the guidance features, such as the cross-level guided high-ratio upsampling from $C4$ to ${\widetilde{C4}}$ in Fig. 1 (b) without complicated iterative procedures as in SAPA.
2) Secondly, for the well-aligned pairs, we propose a paired central difference convolution (PCDC) to inherently capture the local relevance between the query-key features, and then correspondingly construct a parameterized PCDC-Block to flexibly calculate the similarity between every pair. Compared with the non-parametric inner product in Eq. (1), the proposed PCDC-Block can help better ameliorate blurry artifacts, as experimentally validated by comparing ${\widetilde{C4}}$ in Figs. 1 (a) and (b).
3) Thirdly, we propose a fine-grained neighbor selection strategy by selecting the neighbors $\mathcal{N}(i)$ of every HR pixel $i$ based on the bilinearly-upsampled HR version of the LR features $\bm{k}$ and $\bm{x}$. Such a simple yet effective design introduces no extra parameters and can be seamlessly incorporated into our ReSFU framework with ease, which evidently reduces mosaic artifacts even in case of the direct high-ratio upsampling, as presented in Fig. 1 (d).
@@
4) Finally, by correspondingly replacing the three main components in Eq. (1) with our proposed aforementioned three designs, we systematically build ReSFU. Extensive experiments comprehensively demonstrate that compared to baselines, the proposed ReSFU consistently (i) achieves better results in direct high-ratio upsampling on various types of architectures, and (ii) shows superior generality across different segmentation tasks, such as, semantic/instance/panoptic segmentation, and across various types of datasets, including natural and medical images. Moreover, we provide detailed model visualizations and ablation studies to verify the working mechanism underlying ReSFU.
@@
2 Related Work
2.1 Feature Upsampling
@@
2.2 Convolution Operators
To adapt to different application needs, many convolution operators have emerged as replacements for standard convolution in recent years [24]. For instance, based on local binary pattern [25], LBC [26, 27] proposed a set of fixed sparse pre-defined binary convolutional filters, which can significantly save the computational overhead while approximating the capability of traditional convolutional layers. CDC [28, 29, 30, 31] constructed a specific convolution based on central difference maps to capture boundary information. Along this line, PDC [32, 33] further considered the difference between different pixel pairs and SDC [34] introduced both central difference and similarity into the convolution design. They have been effectively applied to vision tasks, such as face anti-spoofing, edge detection, and semantic segmentation. These convolution operators are usually executed on a single input feature. However, in this paper, we aim to novelly design a convolution operation for flexible similarity computation between a pair of content-aligned input features.
2.3 Image Segmentation
In this field of computer vision, image segmentation is a fundamental task in widespread applications, such as medical imaging [35] and autonomous driving [36], which includes semantic segmentation [4], instance segmentation [5], and panoptic segmentation [6]. The goal is to predict semantic labels or delineate individual instances of objects in images, helping provide essential scene understanding.
Against these tasks, various network architectures have been constructed in the past few years. For example, for scene parsing and semantic segmentation of natural images, CNN-based PSPNet [37] and DeepLab-V3+ [38] achieved prominent outcomes by utilizing atrous convolutions and spatial pyramid pooling to effectively capture multi-scale information. Inspired by the successful application of Transformer in computer vision [39, 40], various Transformer-based network architectures have been proposed for semantic segmentation, such as Segmenter [3], SegFormer [1], and SegNeXt [41], which demonstrated impressive performance by leveraging the power of long-range dependencies. For instance segmentation, Mask R-CNN [42] and its variants [43, 44, 45] have been widely used due to the robust performance. For panoptic segmentation, the representative Panoptic FPN [6], UPSNet [46], and EfficientPS [47] sequentially consist of two procedures, i.e., semantic segmentation and instance segmentation. Besides, the U-shape network architectures [35, 48, 49, 50] have been broadly utilized for medical image segmentation. Recently, researchers [51, 52, 53] explored an alternative approach to the traditional pixel-wise classification and proposed to adopt a Transformer-based decoder for mask classification, which shows remarkable performance in different image segmentation tasks.
For these network architectures, the spatial size of intermediate features usually gradually reduces to efficiently encode high-level semantic information, and thus feature upsampling is an indispensable module for the final prediction. In this paper, we aim to design a general feature upsampling framework, which can be easily encapsulated into various structures for more accurate segmentation.
@@ 3 Revisiting Similarity-based Feature Upsampling Pipeline
Given an LR deep feature $\bm{x}\in\mathbb{R}^{hw\times C}$ and an HR guidance feature $\bm{y}\in\mathbb{R}^{HW\times c}$, [^1] For simplicity, we aggregate the spatial dimensions $H\times W$ as $HW$. the existing similarity-based feature upsampling framework SAPA [2, 21] aims to upsample $\bm{x}$ to an HR deep feature $\tilde{\bm{x}}\in\mathbb{R}^{HW\times C}$ under the guidance of $\bm{y}$. Specifically, for each HR pixel $i\in\{1,\ldots,HW\}$, the HR deep feature element $\tilde{\bm{x}}_{i}\in\mathbb{R}^{C}$ is generally estimated by weighting its neighboring LR deep feature elements based on an upsampling kernel $\text{Softmax}(\bm{s}_{i})$ as:
@@
where $\bm{x}_{\mathcal{N}(i)}\in\mathbb{R}^{K^{2}\times C}$ denotes the neighboring LR feature elements of pixel $i$; $|{\mathcal{N}(i)}|=K^{2}$ is the number of neighboring LR pixels with a pre-defined kernel size $K$; $\bm{s}_{i}\in\mathbb{R}^{K^{2}}$ is the similarity scores assigned to the $K^{2}$ neighbors of pixel $i$. In the representative base version of SAPA [2], $\bm{s}_{i}$ is computed as:
$\bm{s}_{i}\triangleq\text{sim}(\bm{q}_{i},\bm{k}_{\mathcal{N}(i)})\triangleq\bm{q}_{i}\bm{k}^{T}_{\mathcal{N}(i)},$ (3)
where $\bm{q}$ and $\bm{k}$ are linear projections of $\bm{y}$ and $\bm{x}$, respectively, with $D$ channels; $\bm{q}_{i}\in\mathbb{R}^{D}$ is the $i$-th feature element of $\bm{q}\in\mathbb{R}^{HW\times D}$; $\bm{k}_{\mathcal{N}(i)}\in\mathbb{R}^{K^{2}\times D}$ represents the $K^{2}$ neighboring feature elements of pixel $i$ in $\bm{k}\in\mathbb{R}^{hw\times D}$; $\text{sim}\left(\cdot,\cdot\right)$ is a general similarity function, designed as the inner product between $\bm{q}_{i}$ and $\bm{k}_{\mathcal{N}(i)}$ in SAPA. We regard $\bm{q}$, $\bm{k}$, and $\bm{x}$ as the HR query, LR key, and LR value features, respectively.
As seen, the entire similarity-based feature upsampling pipeline is strongly associated with the design of three main parts: ❶ query-key feature acquisition for facilitating the computation of similarity $\bm{s}_{i}$; ❷ the similarity function $\text{sim}\left(\cdot,\cdot\right)$; ❸ neighbor selection $\mathcal{N}\left(i\right)$. For better understanding, based on the backbone Segmenter-S with a direct ×$\times 4$ upsampling for the last-layer feature under the guidance of a shallow feature (see Fig. 10), we experimentally visualize SAPA in Fig. 2 (a) and identify the following issues:
@@
2) Inflexible Similarity Computation. The conventional inner product operation in Eq. (3) is non-parametric and lacks flexibility in capturing the relations between the query and key. Directly adopting this inner product form would cause that within a small local region $\mathcal{N}(i)$, the computed $K^{2}$ scores $\bm{q}_{i}\bm{k}^{T}_{\mathcal{N}(i)}$ are close to each other (see Fig. 6 (a)). This would consequently generate an overly smooth upsampling kernel, leading to blurry artifacts in the upsampled $\tilde{\bm{x}}$.

### v2

@@ Abstract
Feature upsampling is a fundamental and indispensable ingredient of almost all current network structures for dense prediciton tasks. Very recently, a popular similarity-based feature upsampling pipeline has been proposed, which utilizes a high-resolution (HR) feature as guidance to help upsample the low-resolution (LR) deep feature based on their local similarity. Albeit achieving promising performance, this pipeline has specific limitations in methodological designs: 1) HR query and LR key features are not well aligned in a controllable manner; 2) the similarity between query-key features is computed based on the fixed inner product form, lacking flexibility; 3) neighbor selection is coarsely operated on LR features, resulting in mosaic artifacts. These shortcomings make the existing methods along this pipeline primarily applicable to hierarchical network architectures with iterative features as guidance and they are not readily extended to a broader range of structures, especially for a direct high-ratio upsampling. Against these issues, we thoroughly refresh this pipeline and meticulously optimize every methodological design. Specifically, we firstly propose an explicitly controllable query-key feature alignment from both semantic-aware and detail-aware perspectives, and then construct a parameterized paired central difference convolution block for flexibly calculating the similarity between the well-aligned query-key features. Besides, we develop a fine-grained neighbor selection strategy on HR features, which is simple yet effective for alleviating mosaic artifacts. Based on these careful designs, we systematically construct a refreshed similarity-based feature upsampling framework named ReSFU. Comprehensive experiments substantiate that our proposed ReSFU is finely applicable to various types of architectures in a direct high-ratio upsampling manner, and consistently achieves satisfactory performance on different applications, including semantic segmentation, medical image segmentation, instance segmentation, panoptic segmentation, object detection, and monocular depth estimation, showing superior generality and ease of deployment beyond the existing upsamplers. Codes are available at https://github.com/zmhhmz/ReSFU.
@@ 1 Introduction
Figure 1: Upper: For the ×$\times 8$ upsampling from $C4$ to $\widetilde{C4}$ in the hierarchical SegFormer [1], (a) SAPA [2] iteratively performs three ×$\times 2$ upsampling processes with intermediate features as guidance; (b) our ReSFU obtains clearer object boundaries in a direct one-step ×$\times 8$ upsampling. Lower: For the ×$\times 4$ upsampling from $\bm{x}$ to $\tilde{\bm{x}}$ under the guidance of the shallow HR feature $\bm{y}$ in the non-hierarchical Segmenter [3], SAPA with (c) the default iterative upsampling incurs mosaic artifacts, which is worse in the (d) direct upsampling, while (e) ReSFU can largely eliminate the artifacts with one-step direct upsampling. All the images are best viewed by zooming in on screen, especially to observe mosaic artifacts.
As a fundamental ingredient in deep network architectures, feature upsampling aims to restore the spatial resolution of low-resolution (LR) features, widely used in dense prediction tasks, such as semantic/instance/panoptic segmentation [4, 5, 6], object detection [7], and depth estimation [8].
@@
Against this issue, a novel similarity-based feature upsampling pipeline SAPA [2, 14] has been proposed, which utilizes an HR feature $\bm{y}$ as guidance to help accomplish the upsampling from the LR feature $\bm{x}$ to the HR one $\tilde{\bm{x}}$, which can be generally formulated as (see Sec. 3 for more details):
@@
$\tilde{\bm{x}}_{i}=\underbracket{\text{Softmax}\left(\text{sim}(\bm{q}_{i},\bm{k}^{T}_{\mathcal{N}(i)})\right)}_{\text{upsampling kernel}}\bm{x}_{\mathcal{N}(i)},$ (1)
where the HR query $\bm{q}$ and the LR key $\bm{k}$ are obtained from $\bm{y}$ and $\bm{x}$, respectively; $\mathcal{N}(i)$ is the neighborhood of the pixel $i$; $\text{sim}(\cdot,\cdot)$ denotes the similarity measure. From Eq. (1), we can easily see that the similarity-based feature upsampling framework mainly consists of three design components: ❶ the acquisition of query-key features $\bm{q}$ and $\bm{k}$; ❷ the similarity measure $\text{sim}(\cdot,\cdot)$ between the query-key pair; ❸ the neighbor selection $\mathcal{N}(i)$. Clearly, for the pipeline, the core essence of generating the high-quality $\tilde{\bm{x}}$ is to systematically optimize these three parts.
@@
By carefully delving into the representative base version of SAPA, we find that each of the aforementioned components in SAPA suffers from unresolved issues. Specifically, ❶ SAPA directly uses the linear projection of the HR guidance feature $\bm{y}$ and the LR deep feature $\bm{x}$ to obtain the query $\bm{q}$ and the key $\bm{k}$, respectively. Considering that $\bm{y}$ is usually shallower than $\bm{x}$, only with a content-agnostic linear projection operation, the generated $\bm{q}$ and $\bm{k}$ are generally not well semantically aligned in a controllable manner. This would adversely interfere with the subsequent similarity calculation; ❷ For $\text{sim}(\cdot,\cdot)$, SAPA simply uses the inner product form, which lacks flexibility and often results in blurry effects in the upsampled features (see $\widetilde{C4}$ in Fig. 1 (a)); ❸ For $\mathcal{N}(i)$, SAPA coarsely selects the neighbors on the LR features $\bm{k}$ and $\bm{x}$, which would incur mosaic artifacts, especially in a direct high-ratio upsampling (see $\tilde{\bm{x}}$ in Figs. 1 (c)(d)).
@@
These three methodological design limitations weaken the potential application of SAPA, making it usually only applicable to iterative ×$\times 2$ upsampling structures with hierarchical guidances by default but less suitable for direct high-ratio upsampling. For instance, for the hierarchical architecture SegFormer [1] with four levels of features as shown in Fig. 1 (a), to accomplish the ×$\times 8$ upsampling from $C4$ to $\widetilde{C4}$, SAPA, as well as other existing feature upsampling methods [9, 15, 16], performs three consecutive ×$\times 2$ upsampling processes with $C3,C2$, and $C1$ as guidance features, respectively. Such an iterative pattern definitely leads to the laborious deployment of more upsampling modules. If incorporating them in a direct high-ratio upsampling manner like Fig. 1 (b), the performance generally drops (see Sec. 5.5). Besides, we pioneer the application of existing upsampling methods on non-hierarchical structures, such as Segmenter [3]. As shown in Figs. 1 (c) and (d), for the $\times$4 high-ratio upsampling, whether inserting SAPA in the iterative upsampling manner by default or in the direct high-ratio upsampling manner causes different degrees of mosaic artifacts. These issues generally exist in most of the current feature upsampling methods (see Sec. 5.2).
Faced with this tricky situation, a crucial question inevitably arises: is it possible to build an architecture-agnostic feature upsampling framework that is applicable to different types of network architectures and can always achieve satisfactory effects in direct high-ratio upsampling without complicated iterative processes?
To answer this question, in this paper, we meticulously refresh the pipeline in Eq. (1) and carefully optimize the involved three main components one by one. ❶ Firstly, we propose an explicitly controllable algorithm for the effective query-key alignment of $\bm{q}$ and $\bm{k}$. Concretely, we utilize the guided filter (GF) [17] and Gaussian smoothing to accomplish semantic-aware mutual-alignment and detail-aware self-alignment, respectively. Such careful query-key alignment designs facilitate more accurate similarity computation and make it possible to directly execute the one-step high-ratio upsampling without complicated iterative procedures as in SAPA (see Figs. 1 (b)(e)). ❷ Secondly, for $\text{sim}(\cdot,\cdot)$, we propose a paired central difference convolution (PCDC) to inherently capture the local relevance between the query-key features, and then correspondingly construct a parameterized PCDC-Block to flexibly calculate the similarity between every well-aligned pair. Compared with the non-parametric inner product, the proposed PCDC-Block can help better ameliorate blurry artifacts, as experimentally validated by comparing ${\widetilde{C4}}$ in Figs. 1 (a) and (b). ❸ Thirdly, we propose a fine-grained neighbor selection strategy by selecting the neighbors $\mathcal{N}(i)$ of every HR pixel $i$ based on the bilinearly-upsampled HR version of the LR features $\bm{k}$ and $\bm{x}$. Such a simple yet effective design introduces no extra parameters, which evidently reduces mosaic artifacts even in case of the direct high-ratio upsampling, as presented in Fig. 1 (e). Based on the proposed formulation in Eq. (1) and these three aforementioned methodological designs, we systematically construct a Refreshed Similarity-based Feature Upsampling framework, called ReSFU. In summary, our main contributions are three-fold:
1) To the best of our knowledge, we are the first to provide the mathematical formulation of the similarity-based feature upsampling framework and reveal the underlying optimization essence (see Eq. (1)). Motivated by the essential understanding, we comprehensively optimize the three components. Correspondingly, we propose the explicitly controllable query-key alignment strategy to acquire the query-key features, PCDC-Block for similarity calculation $\text{sim}\left(\cdot,\cdot\right)$, and fine-grained neighbor selection $\mathcal{N}(i)$. These three essential methodological designs harmoniously form an inseparable whole as our proposed ReSFU.
2) Attributed to our meticulous methodological designs, the proposed ReSFU is equipped with multiple application advantages. Not only can it be inserted multiple steps iteratively like Figs. 1 (a)(c) adopted by previous upsampling methods, but more meaningfully, even using the more difficult one-step high-ratio upsampling manner like Figs. 1 (b)(e), our proposed ReSFU can still obtain superior performance while producing visually satisfactory effects (see Secs. 5.2 and 5.5). Moreover, with the default configuration of direct high-ratio upsampling, our ReSFU can be easily and widely deployed to various network types, including hierarchical and non-hierarchical architectures, without requiring iterative intermediate features as guidance, which shows excellent generality and universality (see Fig. 8).
@@
3) Extensive experiments comprehensively demonstrate that compared to baselines, the proposed ReSFU consistently (i) achieves better results in a direct high-ratio upsampling on various types of architectures, and (ii) shows superior generality across different tasks, such as, semantic/instance/panoptic segmentation, object detection, and monocular depth estimation, and across various types of datasets, including natural and medical images. Moreover, we provide detailed model visualizations and ablation studies to verify the working mechanism of every component in the proposed ReSFU.
@@
2 Related Work on Feature Upsampling
@@
Please refer to the supplementary material (SM) for more extensive related works, such as the convolution operators and image segmentation methods.
@@ 3 Revisiting Similarity-based Feature Upsampling Pipeline
Given an LR deep feature $\bm{x}\in\mathbb{R}^{hw\times C}$ and an HR guidance feature $\bm{y}\in\mathbb{R}^{HW\times c}$, [^1] For simplicity, we aggregate the spatial dimensions $H\times W$ as $HW$. the similarity-based feature upsampling framework SAPA [2, 14] aims to upsample $\bm{x}$ to an HR deep feature $\tilde{\bm{x}}\in\mathbb{R}^{HW\times C}$ under the guidance of $\bm{y}$. Specifically, for each HR pixel $i\in\{1,\ldots,HW\}$, the HR deep feature element $\tilde{\bm{x}}_{i}\in\mathbb{R}^{C}$ is generally estimated by weighting its neighboring LR deep feature elements based on an upsampling kernel $\text{Softmax}(\bm{s}_{i})$ as:
@@
where $\bm{x}_{\mathcal{N}(i)}\in\mathbb{R}^{K^{2}\times C}$ denotes the neighboring LR feature elements of pixel $i$; $|{\mathcal{N}(i)}|=K^{2}$ is the number of neighboring LR pixels with a pre-defined kernel size $K$; $\bm{s}_{i}\in\mathbb{R}^{K^{2}}$ is the similarity scores assigned to the $K^{2}$ neighbors of pixel $i$, which is generally computed as:
$\bm{s}_{i}=\text{sim}(\bm{q}_{i},\bm{k}_{\mathcal{N}(i)}),$ (3)
where $\bm{q}$ and $\bm{k}$ are generated from $\bm{y}$ and $\bm{x}$, respectively, with $D$ channels; $\bm{q}_{i}\in\mathbb{R}^{D}$ is the $i$-th feature element of $\bm{q}\in\mathbb{R}^{HW\times D}$; $\bm{k}_{\mathcal{N}(i)}\in\mathbb{R}^{K^{2}\times D}$ represents the $K^{2}$ neighboring feature elements of pixel $i$ in $\bm{k}\in\mathbb{R}^{hw\times D}$; $\text{sim}\left(\cdot,\cdot\right)$ is a general similarity function. We regard $\bm{q}$, $\bm{k}$, and $\bm{x}$ as the HR query, LR key, and LR value features, respectively.
As seen, the entire similarity-based feature upsampling pipeline is strongly associated with the design of three main parts: ❶ query-key feature acquisition for facilitating the computation of similarity $\bm{s}_{i}$; ❷ the similarity function $\text{sim}\left(\cdot,\cdot\right)$; ❸ neighbor selection $\mathcal{N}\left(i\right)$.
For the representative base version of SAPA [2], the query $\bm{q}$ and the key $\bm{k}$ are obtained by taking linear projections of $\bm{y}$ and $\bm{x}$, denoted as $\text{Proj}_{\bm{\theta}_{\bm{q}}}(\bm{y})$ and $\text{Proj}_{\bm{\theta}_{\bm{k}}}(\bm{x})$, respectively, where $\bm{\theta}_{\bm{q}}$ and $\bm{\theta}_{\bm{k}}$ are network parameters. The similarity measure is implemented via a simple inner product operation, i.e., $\text{sim}(\bm{q}_{i},\bm{k}_{\mathcal{N}(i)})=\bm{q}_{i}\bm{k}^{T}_{\mathcal{N}(i)}$. The neighbors $\mathcal{N}(i)$ of pixel $i$ are selected coarsely on the LR features. For better understanding, based on the backbone Segmenter-S with a direct ×$\times 4$ upsampling for the last-layer feature under the guidance of a shallow feature (see Fig. 8 (b)), we experimentally visualize SAPA in Fig. 2 (a) and identify the following issues:
@@
2) Inflexible Similarity Computation. The conventional inner product operation is non-parametric and lacks flexibility in capturing the relations between the query and key. Directly adopting this inner product form would cause that within a small local region $\mathcal{N}(i)$, the computed $K^{2}$ scores $\bm{q}_{i}\bm{k}^{T}_{\mathcal{N}(i)}$ are close to each other (see Fig. 6 (a)). This would consequently generate an overly smooth upsampling kernel, leading to blurry artifacts in the upsampled $\tilde{\bm{x}}$.